OpenStudy (anonymous):

How to solve for x? 2x-2/4 (whole thing is divided by 4) = 2x-4/6 (whole thing is divided by 6)

OpenStudy (nubeer):

2 (x-1)/4 = 2 (x-2) / 6 (x-1)/2 = (x-2) /3 cross multiply 3x-3 = 2x-4 x = -1

OpenStudy (anonymous):

Thank you very much! I have a teacher who just gave us these questions but didn't teach us any rules. He also doesn't use a textbook so I am going to be looking for the best materials to learn this stuff. Any suggestions? I am in seventh grade.

OpenStudy (nubeer):

hmm well its hard if u don't even have text book.. i personally never searched much online so don't know about online material.. well i explained it to u.. first you have to see if u can take any factor common and take it out if possible.. u can solve by cross multiplication or u can also solve by taking LCM ..

OpenStudy (anonymous):

Thanks very much.

OpenStudy (nubeer):

No problem :) its just multiplication and algebra.. tryto write every step and be careful with signs.. the rest would be easy :)